OpenStudy (anonymous):

x^2 - 8x + 16 = 0 **For each equation , evaluate the discriminant & determine the nature of the solutions. Then solve each equation using the quadratic formula to verify the nature of the roots

OpenStudy (anonymous):

\[a x^2+b x+c=0\] \[discriminant~ D=b^2-4 ac\] if D>0,roots are real \[if D \ge 0~ and~ a~ perfect \square,roots~are~real~rational~and~different.\] if D=0,roots are real and equal. if D<0,roots are imaginary.